# A Hierarchical Framework for Measuring Scientific Paper Innovation via Large Language Models

### v1

@@ V-E3 HSPIM Framework Comparison
The bold values represent the best results.
TABLE VI: Semantic textual similarity between LLM-generated scoring reasons and actual peer review comments
@@
[…]
@@
The bold values represent the best results.
V-F Semantic Textual Similarity
Table VI reports the semantic similarity between scoring rationales generated by LLMs and actual peer review comments. The experiment compares SEA-E, SSPIM, and HSPIM in terms of how closely their generated evaluation texts align with real reviews. In SSPIM and HSPIM, LLMs are prompted to output not only the “novelty_score” and “confidence_score” but also a “reason” to explain the novelty score. To measure semantic similarity, we use cosine similarity and BERTScore [48]. Cosine similarity is computed with sentence embeddings from the gte-large-en-v1.5 model [51]. BERTScore uses the longformer-base-4096 [52] model to support long inputs. HSPIMjoint achieves higher scores than SEA-E and SSPIM, with an average Cosine Similarity improvement of 27.3% and 7.4%, and a BERTScore improvement of 3.6% and 0.7%, respectively. Compared to SEA-E and SSPIM, HSPIM generates scoring rationales that are semantically closer to real peer review comments. Therefore, our HSPIM framework can not only assess innovation quantitatively but also explain it qualitatively.
TABLE VII: RMSE comparison of different LLM-based HSPIMjoint
@@ V-F Semantic Textual Similarity
The bold values represent the best results.
V-G LLM-based HSPIM Comparison
Table VII compares the RMSE results of HSPIMjoint using different LLMs. In addition to DeepSeek-V3 and GPT-4o Mini, we also deploy three small open-source LLMs: LLaMA-3-8B [53], Mistral-7B [54], and Qwen2.5-7B [55]. We see that HSPIM performs well on small-scale LLMs, showing its potential to support innovation assessment with lightweight models. Notably, Mistral-7B outperforms other LLMs on three out of four datasets.
(a) ICLR-2017
(b) CoNLL-2016
(c) ACL-2017
(d) COLING-2020
[…]
(a) ACL-2017
(b) CoNLL-2016
(c) ICLR-2017
(d) COLING-2020
@@ V-G LLM-based HSPIM Comparison
The bold values represent the best results.
VI Experimental Analysis
VI-A Ablation Study
VI-A1 Ablation Study on Confidence Score
Table VIII compares the performance with (w/) and without (w/o) using confidence scores. The table reports RMSE results for SSPIM and HSPIMjoint on DeepSeek-V3. Using confidence scores to do weighted novelty scoring helps to reduce RMSE in both SSPIM and HSPIM frameworks. This result shows that adding confidence scores is helpful to connect local novelty evaluations to full-paper innovation.
TABLE IX: Effectiveness of two-layer question structure on HSPIM framework
@@ VI-A1 Ablation Study on Confidence Score
The bold values represent the best results.
TABLE X: RMSE comparison of representative section combinations on HSPIM
@@
The bold values represent the best results.
VI-A2 Ablation Study on Two-Layer Question Structure
Table IX presents the results of the ablation study on two-layer question structure under DeepSeek-HSPIM. We can see that combining specific and common questions helps to reduce RMSE by 13.90% compared to using only specific questions, and by 18.43% compared to using only common questions. These results confirm the effectiveness of the two-layer question structure.
VI-B Analysis of Section-Based Scoring
VI-B1 Effect of Section Selection on Innovation Scoring
Fig. 6 compares the RMSE of using individual sections (scored by LLM one-step prompting), SSPIM, and HSPIMjoint across four datasets. We observe that HSPIMjoint consistently achieves the lowest RMSE on ICLR-2017, CoNLL-2016, and COLING-2020, and also outperforms five out of nine single-section baselines on ACL-2017. The results highlight the effectiveness of section-based weighted novelty scoring (from individual sections to SSPIM and HSPIM). In addition, HSPIM provides more stable RMSE performance than using single sections.
VI-B2 Scoring Bias Across Sections
VI-B3 Modeling Section Combination
@@
Table X shows the RMSE results for different section combining strategies using GPT-HSPIMjoint. We group these strategies into five categories based on the functional relationship between the merged sections. For example, the combination Approach + Experiments merges the method and experiment content to capture their complementary information. All Sections, Independent denotes the original setup where no sections are merged. We observe that some section combinations outperform the original setting on certain datasets. Compared to the original setup, HSPIM with combined sections shows performance gains in several cases, such as Approach + Experiments, Introduction + Discussion + Conclusion, Introduction + Related Work + Approach + Experiments, and Experiments + Analysis + Discussion + Conclusion. Hence, combining sections is a potential improvement path for the HSPIM framework.
TABLE XI: Effect of applying neutralization to abstract and introduction
@@
VI-C Sensitivity Analysis on Self-Praising Expressions
Scientific paper sections such as Abstract and Introduction often include self-praising expressions, such as “significantly better” or “a novel model.” To analyze how such expressions affect the innovation scoring in HSPIM framework, we conduct a sensitivity analysis on DeepSeek-HSPIMjoint and show the results in Table XI. Note that this analysis only focus on testing the robustness of HSPIM to self-praising expressions. It does not assume that the replaced expressions are necessarily exaggerated or unreasonable. Specifically, we use zero-shot LLMs to detect self-praising expressions in the Abstract and Introduction of each paper. Then, LLMs rewrite these expressions in a neutral tone. For example, LLMs may change “significantly better” to “better” and “a novel model” to “a new model.”
In Table XI, +Neutralize rewrites self-praising expressions in the Abstract and Introduction into more neutral statements using DeepSeek. We observe that +Neutralize notably reduces the average innovation score for HSPIM with only Abstract + Introduction. This suggests that zero-shot LLMs tend to assign higher scores when the text contains subjective or emphatic language. On the other hand, HSPIM with all section types (noted as HSPIM (All Sections)) is more robust, showing smaller changes in both innovation scores and RMSE. HSPIM combines text chunks from all section types to improve the stability of the innovation scoring.
VI-D Comparison of Prompt Optimization Methods
@@
TABLE XII: Computational cost of different LLMs

### v2

@@ 5.5 Main Results
5.5.3 HSPIM Framework Comparison
Table 5 shows the performance comparison of three framework variants: section-based scientific paper innovation measurement (SSPIM), HSPIM using the initial prompt combination (naive HSPIM), and HSPIM with GA-based joint optimization strategy (HSPIMjoint). We report results on both the training set and the test set. We observe a consistent decrease in RMSE from SSPIM to naive HSPIM, and further to HSPIMjoint, which highlights the importance of each component. Specifically, naive HSPIM outperforms SSPIM, which shows the effectiveness of QA augmentation. HSPIMjoint improves naive HSPIM, which confirms the benefit of multi-prompt optimization.
Table 6: Performance of GPT-HSPIM+ with different norms and optimization strategies across datasets.
Norm Model ICLR-2017 CoNLL-2016 ACL-2017 COLING-2020 Average
RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE
$L_{1}$ HSPIM+joint 0.7695 0.6090 0.3992 0.3947 0.9461 0.7865 0.4604 0.3806 0.6438 0.5427
HSPIM+two-step 0.7904 0.6372 0.4279 0.4128 0.9386 0.7801 0.4529 0.3870 0.6525 0.5543
HSPIM+pruning=3 0.3944 0.2667 0.2873 0.2870 0.8440 0.7262 0.2873 0.2870 0.4533 0.3917
$L_{2}$ HSPIM+joint 0.7693 0.6078 0.4038 0.3993 0.9425 0.7848 0.4621 0.3827 0.6444 0.5437
HSPIM+two-step 0.7893 0.6349 0.4318 0.4169 0.9348 0.7781 0.4533 0.3873 0.6523 0.5543
HSPIM+pruning=3 0.3970 0.2714 0.2986 0.2981 0.8409 0.7246 0.2986 0.2981 0.4588 0.3980
$L_{\infty}$ HSPIM+joint 0.8142 0.5864 0.6336 0.6254 0.7978 0.7047 0.5833 0.4795 0.7072 0.5990
HSPIM+two-step 0.7845 0.5655 0.6473 0.6389 0.7854 0.6899 0.5330 0.4569 0.6875 0.5878
HSPIM+pruning=3 0.5217 0.3667 0.6473 0.6389 0.7676 0.6786 0.6473 0.6389 0.6460 0.5808
5.5.4 HSPIM+ Results
Following Definition 3.2, HSPIM+ measures innovation as a norm-based combination of novelty, contribution, and feasibility. As shown in Table 6, the choice of aggregation norm ($L_{1}$, $_{2}, LL_{\infty}$) has a notable effect on performance.
The results in Table 6 show that the $L_{1}$ and $L_{2}$ norms yield similar performance across all datasets and optimization strategies. Notably, the $L_{\infty}$ norm achieves the lowest errors on the ACL-2017 dataset, outperforming both $L_{1}$ and $L_{2}$. For all norm choices, the pruning strategy achieves the lowest RMSE and MAE, indicating its effectiveness.
Compared to the original HSPIM framework (Table 3), HSPIM+ achieves a lower best average RMSE and MAE (0.4533 and 0.3917) than HSPIM (0.5915 and 0.4681). This improvement is particularly clear on ACL-2017, where HSPIM+ outperforms zero-shot LLM baselines and matches the performance of supervised models. These results demonstrate that modeling innovation as a composite of novelty, contribution, and feasibility through norm-based aggregation is both theoretically justified and empirically validated.
Table 7: Performance comparison on a subset of the PMC Article Dataset and the arXiv physics.ins-det corpus. Bold highlights the best results and underline highlights the second best results.
Datasets PMC arXiv
Metrics RMSE MAE RMSE MAE
Supervised Models
CNN 1.5481 1.3022 1.4870 1.2244
RNN 1.0187 0.7878 1.1123 0.8823
Transformer 0.8884 0.6814 0.9380 0.7540
Zero-Shot LLMs
GPT-4o mini 3.1006 2.9105 1.5711 1.3542
DeepSeek-V3 3.2957 3.1779 1.2333 1.0221
GPT-HSPIM 0.8824 0.6853 1.0371 0.8256
DeepSeek-HSPIM 0.8956 0.6755 0.9424 0.7554
5.5.5 Domain Generalization Experiments
To evaluate the domain generalization performance of the HSPIM framework, we conduct experiments on subsets of the PubMed Central (PMC) Article Datasets and the arXiv physics.ins-det (Instrumentation and Detectors) collection. Since both datasets lack peer review scores, following previous studies, we use citation counts as a proxy for innovation scores [5]. We then map citation counts to ground-truth innovation scores via a nonlinear transformation that approximates a normal distribution, so that the resulting score distribution closely fits the ground-truth innovation score distribution observed in Section 6.5. For PMC, we select 95 biochemical articles published in 2004 from the journal Bioinformatics to ensure uniformity in publication venue, year, and domain. For arXiv, we use the 2018 corpus of physics.ins-det (Instrumentation and Detectors), comprising 276 papers on detector design, readout and DAQ, performance evaluation, simulation, and AI-assisted control.
As shown in Table 7, HSPIM achieves the best RMSE and MAE on PMC and the second-best scores on arXiv, and it outperforms zero-shot LLM baselines on both datasets. These results indicate strong cross-domain performance and support the general applicability of HSPIM.
Table 8: Semantic textual similarity between LLM-generated scoring reasons and actual peer review comments. The bold values represent the best results.
@@ 5.5.5 Domain Generalization Experiments
$_{\text{joint}} SEA-E SSPIM HSPIM$$_{\text{joint}}$
@@
5.6 Semantic Textual Similarity
Table 8 reports the semantic similarity between scoring rationales generated by LLMs and actual peer review comments. The experiment compares SEA-E, SSPIM, and HSPIM in terms of how closely their generated evaluation texts align with real reviews. In SSPIM and HSPIM, LLMs are prompted to output not only the “novelty_score” and “confidence_score” but also a “reason” to explain the novelty score. To measure semantic similarity, we use cosine similarity and BERTScore [41]. Cosine similarity is computed with sentence embeddings from the gte-large-en-v1.5 model [44]. BERTScore uses the longformer-base-4096 [45] model to support long inputs. HSPIMjoint achieves higher scores than SEA-E and SSPIM, with an average Cosine Similarity improvement of 27.3% and 7.4%, and a BERTScore improvement of 3.6% and 0.7%, respectively. Compared to SEA-E and SSPIM, HSPIM generates scoring rationales that are semantically closer to real peer review comments. Therefore, our HSPIM framework can not only assess innovation quantitatively but also explain it qualitatively.
Table 9: RMSE comparison of different LLM-based HSPIMjoint. The bold values represent the best results.
@@ 5.6 Semantic Textual Similarity
5.7 LLM-based HSPIM Comparison
Table 9 compares the RMSE results of HSPIMjoint using different LLMs. In addition to DeepSeek-V3 and GPT-4o Mini, we also deploy three small open-source LLMs: LLaMA-3-8B [46], Mistral-7B [47], and Qwen2.5-7B [48]. We see that HSPIM performs well on small-scale LLMs, showing its potential to support innovation assessment with lightweight models. Notably, Mistral-7B outperforms other LLMs on three out of four datasets.
@@ 5.7 LLM-based HSPIM Comparison
Table 11: Effectiveness of two-layer question structure on HSPIM framework. The bold values represent the best results.
@@
6 Experimental Analysis
6.1 Ablation Study
In this section, we conduct ablation study on confidence score and two-layer question structure.
6.1.1 Ablation Study on Confidence Score
Table 10 compares the performance with (w/) and without (w/o) using confidence scores. The table reports RMSE results for SSPIM and HSPIMjoint on DeepSeek-V3. Using confidence scores to do weighted innovation scoring helps to reduce RMSE in both SSPIM and HSPIM frameworks. This result shows that adding confidence scores is helpful to connect local novelty evaluations to full-paper innovation.
6.1.2 Ablation Study on Two-Layer Question Structure
Table 11 presents the results of the ablation study on two-layer question structure under DeepSeek-HSPIM. We can see that combining specific and common questions helps to reduce RMSE by 13.90% compared to using only specific questions, and by 18.43% compared to using only common questions. These results confirm the effectiveness of the two-layer question structure.
(a) ICLR-2017
(b) CoNLL-2016
(c) ACL-2017
(d) COLING-2020
[…]
(a) ACL-2017
(b) CoNLL-2016
(c) ICLR-2017
(d) COLING-2020
6.2 Analysis of Section-Based Scoring
In this section, we conduct analyses of section-based scoring.
6.2.1 Effect of Section Selection on Innovation Scoring
Fig. 7 compares the RMSE of using individual sections (scored by LLM one-step prompting), SSPIM, and HSPIMjoint across four datasets. We observe that HSPIMjoint consistently achieves the lowest RMSE on ICLR-2017, CoNLL-2016, and COLING-2020, and also outperforms five out of nine single-section baselines on ACL-2017. The results highlight the effectiveness of section-based weighted innovation scoring (from individual sections to SSPIM and HSPIM). In addition, HSPIM provides more stable RMSE performance than using single sections.
6.2.2 Scoring Bias Across Sections
Table 12: RMSE comparison of representative section combinations on HSPIM. The bold values represent the best results.
@@ 6.2.2 Scoring Bias Across Sections
6.2.3 Modeling Section Combination
@@
Table 12 shows the RMSE results for different section combining strategies using GPT-HSPIMjoint. We group these strategies into five categories based on the functional relationship between the merged sections. For example, the combination Approach + Experiments merges the method and experiment content to capture their complementary information. All Sections, Independent denotes the original setup where no sections are merged. We observe that some section combinations outperform the original setting on certain datasets. Compared to the original setup, HSPIM with combined sections shows performance gains in several cases, such as Approach + Experiments, Introduction + Discussion + Conclusion, Introduction + Related Work + Approach + Experiments, and Experiments + Analysis + Discussion + Conclusion. Hence, combining sections is a potential improvement path for the HSPIM framework.
6.3 Sensitivity Analysis
In this section, we conduct sensitivity analyses of HSPIM with respect to LLM temperature, prompt instructions for critical scoring, and self-praising expressions in scientific papers.
Table 13: Performance of GPT-HSPIMjoint under different LLM temperature settings.
Dataset LLM Temperature
0 0.3 0.7 1.0
RMSE MAE RMSE MAE RMSE MAE RMSE MAE
ICLR-2017 0.8474 0.7160 0.8499 0.6399 0.8399 0.6359 0.8532 0.6463
CoNLL-2016 0.1973 0.1736 0.0663 0.0593 0.1321 0.1197 0.1587 0.1831
ACL-2017 1.1205 0.8897 1.1489 0.9071 1.1432 0.9629 1.1395 0.9267
COLING-2020 0.5075 0.4129 0.4872 0.3901 0.4844 0.3875 0.4593 0.3666
6.3.1 Sensitivity Analysis on LLM Temperature
We analyze the effect of the LLM temperature parameter on the performance of GPT-HSPIMjoint. We analyze the effect of the LLM temperature parameter on the performance of GPT-HSPIMjoint. In all other experiments, we set the temperature to 0 to reduce randomness. To increase diversity, we set the temperature to higher values such as $1$.
As shown in Table 13, varying the temperature from 0 to $1.0$ has only a limited impact on RMSE and MAE across all datasets. The results show that increasing the temperature does not significantly change model predictions or reduce conservative scoring. HSPIM framework remains robust to this parameter.
Table 14: Effect of adding the critical scoring instruction in the scoring prompt on GPT-HSPIMjoint.
Dataset Setting RMSE MAE
ICLR-2017 Original 0.8474 0.7160
+Critical Scoring 0.7939 0.6752
Change (%) -6.31% -5.70%
CoNLL-2016 Original 0.1973 0.1736
+Critical Scoring 0.1557 0.1464
Change (%) -21.08% -15.67%
ACL-2017 Original 1.1205 0.8897
+Critical Scoring 1.1614 0.9436
Change (%) +3.65% +6.06%
COLING-2020 Original 0.5075 0.4129
+Critical Scoring 0.4983 0.4049
Change (%) -1.81% -1.94%
6.3.2 Sensitivity Analysis on Critical Scoring instruction
Table 14 shows the performance of GPT-HSPIMjoint with and without an additional instruction in the scoring prompt. Specially, we add the instruction “Be decisive in your scoring and give higher or lower scores when there is clear evidence, rather than always choosing the middle range” to the scoring prompt to encourage critical evaluation. This experiment tests whether HSPIM is sensitive to the scoring instruction. The analysis also aims to find out if this instruction can reduce conservative scores and improve results. We pay attention to the ACL-2017 dataset to see if this change helps HSPIM perform better.
The results show that adding the critical scoring instruction lowers RMSE and MAE on most datasets, suggesting HSPIM is sensitive to prompt wording. However, this approach does not improve results on ACL-2017. Such an approach can help address the narrow distribution of predicted scores shown in Fig. 10, but more effective strategies are still required. We plan to explore better solutions in future work.
Table 15: Effect of applying neutralization to abstract and introduction.
@@ 6.3.2 Sensitivity Analysis on Critical Scoring instruction
6.3.3 Sensitivity Analysis on Self-Praising Expressions
Scientific paper sections such as Abstract and Introduction often include self-praising expressions, such as “significantly better” or “a novel model.” To analyze how such expressions affect the innovation scoring in HSPIM framework, we conduct a sensitivity analysis on DeepSeek-HSPIMjoint and show the results in Table 15. Note that this analysis only focus on testing the robustness of HSPIM to self-praising expressions. It does not assume that the replaced expressions are necessarily exaggerated or unreasonable. Specifically, we use zero-shot LLMs to detect self-praising expressions in the Abstract and Introduction of each paper. Then, LLMs rewrite these expressions in a neutral tone. For example, LLMs may change “significantly better” to “better” and “a novel model” to “a new model.”
In Table 15, +Neutralize rewrites self-praising expressions in the Abstract and Introduction into more neutral statements using DeepSeek. We observe that +Neutralize notably reduces the average innovation score for HSPIM with only Abstract + Introduction. This suggests that zero-shot LLMs tend to assign higher scores when the text contains subjective or emphatic language. On the other hand, HSPIM with all section types (noted as HSPIM (All Sections)) is more robust, showing smaller changes in both innovation scores and RMSE. HSPIM combines text chunks from all section types to improve the stability of the innovation scoring.
6.4 Comparison of Prompt Optimization Methods
@@
(a) ICLR-2017
(b) CoNLL-2016
(c) ACL-2017
(d) COLING-2020
Figure 10: Comparison of ground-truth and HSPIMjoint predicted innovation score distributions on four datasets. Predicted scores are more concentrated around the mean.
6.5 Analysis of Innovation Score Distributions
Fig. 10 compares the score distributions of ground-truth and HSPIMjoint predictions on four datasets. Across all datasets, ground-truth scores exhibit a wider spread, including a noticeable proportion of both high and low (tail) values. In contrast, HSPIMjoint predictions are more concentrated and have a lower mean and variance. For example, on ACL-2017, the predicted mean and variance are 3.48 and 0.01, while the ground-truth mean and variance are 3.97 and 0.57. In HSPIM, LLMs rarely give extreme scores. The mismatch in distribution helps explain performance drops on datasets with more diverse or higher ground-truth scores.
Table 16: Misclassification handling and label-noise sensitivity for GPT-HSPIMjoint. Noise is injected by randomly converting $p\%$ of section labels to unmatched. $\Delta_{\text{orig}}$: change vs Original; $\Delta_{\text{handle}}$: change vs With misclassification handling.
Dataset Setting RMSE MAE
ICLR-2017 Original 0.8283 0.6966
With misclassification handling 0.8248 ($\Delta_{\text{orig}}$ $0.42\%$) 0.6951 ($\Delta_{\text{orig}}$ $0.22\%$)
+ Noise (unmatched 5%) 0.8376 ($\Delta_{\text{handle}}$ $1.56\%$) 0.7064 ($\Delta_{\text{handle}}$ $1.63\%$)
+ Noise (unmatched 10%) 0.8268 ($\Delta_{\text{handle}}$ $0.25\%$) 0.6972 ($\Delta_{\text{handle}}$ $0.30\%$)
+ Noise (unmatched 20%) 0.8430 ($\Delta_{\text{handle}}$ $2.20\%$) 0.7079 ($\Delta_{\text{handle}}$ $1.85\%$)
CoNLL-2016 Original 0.0638 0.0560
With misclassification handling 0.0798 ($\Delta_{\text{orig}}$ $25.08\%$) 0.0659 ($\Delta_{\text{orig}}$ $17.68\%$)
+ Noise (unmatched 5%) 0.0798 ($\Delta_{\text{handle}}$ $25.08\%$) 0.0659 ($\Delta_{\text{handle}}$ $17.68\%$)
+ Noise (unmatched 10%) 0.0798 ($\Delta_{\text{handle}}$ $25.08\%$) 0.0659 ($\Delta_{\text{handle}}$ $17.68\%$)
+ Noise (unmatched 20%) 0.0798 ($\Delta_{\text{handle}}$ $25.08\%$) 0.0659 ($\Delta_{\text{handle}}$ $17.68\%$)
ACL-2017 Original 1.0887 0.8572
With misclassification handling 1.1491 ($\Delta_{\text{orig}}$ $5.55\%$) 0.9284 ($\Delta_{\text{orig}}$ $8.31\%$)
+ Noise (unmatched 5%) 1.1526 ($\Delta_{\text{handle}}$ $0.31\%$) 0.9270 ($\Delta_{\text{handle}}$ $0.15\%$)
+ Noise (unmatched 10%) 1.1651 ($\Delta_{\text{handle}}$ $1.39\%$) 0.9360 ($\Delta_{\text{handle}}$ $0.82\%$)
+ Noise (unmatched 20%) 1.1566 ($\Delta_{\text{handle}}$ $0.66\%$) 0.9248 ($\Delta_{\text{handle}}$ $0.39\%$)
COLING-2020 Original 0.5202 0.3933
With misclassification handling 0.4806 ($\Delta_{\text{orig}}$ $7.61\%$) 0.3874 ($\Delta_{\text{orig}}$ $1.50\%$)
+ Noise (unmatched 5%) 0.4758 ($\Delta_{\text{handle}}$ $1.00\%$) 0.3846 ($\Delta_{\text{handle}}$ $0.74\%$)
+ Noise (unmatched 10%) 0.4722 ($\Delta_{\text{handle}}$ $1.75\%$) 0.3815 ($\Delta_{\text{handle}}$ $1.53\%$)
+ Noise (unmatched 20%) 0.4810 ($\Delta_{\text{handle}}$ $0.09\%$) 0.3835 ($\Delta_{\text{handle}}$ $1.02\%$)
6.6 Impact of Section Misclassification
To address the risk of inaccurate section assignment for ambiguous or unconventional section content, we introduce a revised classification strategy, referred to as With revised classification strategy in Table 16. Previously, the LLM was required to assign every section to the closest type in a fixed list, even when the fit was poor. In this revised setting, HSPIM first attempts to match each section to the predefined types. If a section cannot be confidently classified, it is labeled as “unmatched” and, in the QA augmentation step of HSPIM, only uses the common question prompt without applying any section-specific questions. This design allows HSPIM to more flexibly handle ambiguous or unconventional section content. As shown in Table 16, this strategy maintains stable performance on most datasets. In addition, we test robustness by randomly converting 5%, 10%, and 20% of section labels to unmatched and re-running QA and scoring. The results show that RMSE and MAE change only slightly, confirming resilience to occasional section misclassification.
Table 17: Computational cost of different LLMs.